OpenStudy (anonymous):

IM SOO HAPPY I THINK I UNDERSTAND THIS!!!Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8 (-2, 0) no solution (3, 4) infinitely many solutions the answer is (-2,0) right?

OpenStudy (radar):

Just check it, substitute the -2 for x and 0 for and see if it works. I personally don't think it will.

OpenStudy (anonymous):

i graphed it

OpenStudy (campbell_st):

rewrite the lines lin point, slope form subtract 2x to both sides add x to both sides -2y = - 1/2 x + 4 4y = x - 8 divide by -2 divide by 4 y = -1/4x -2 y = -1/4x - 2 what can you make of this...

OpenStudy (anonymous):

uuuh ...?

OpenStudy (anonymous):

am i right?

OpenStudy (radar):

(1/2)(-2) - 2 (0) = 4 ??????????

OpenStudy (radar):

Does -1 - 0 = 4 ?

OpenStudy (anonymous):

no...

OpenStudy (radar):

That is what you proposed solution results for the first equation.

OpenStudy (anonymous):

look i graphed it and they intercepted at (-2,0) and that is a choice

OpenStudy (radar):

Well I guess we can't argue with that! If they intersect then that is the answer. Are you sure you graphed them properly?

OpenStudy (anonymous):

yes i checked them

OpenStudy (radar):

Did they actually intersect or were they the same line (coincidence)

OpenStudy (anonymous):

no trhey intersected

OpenStudy (radar):

From what I can tell they are two lines with the same slope intercept equation. y=(1/4)x - 2

OpenStudy (radar):

They are the same line and the are coincidence, with infinite points of coincidence with infinitely number of solutions.

OpenStudy (anonymous):

no i mean they did exactly intercect at -2,0

OpenStudy (radar):

Well let me double check......the first equation is (1/2)x -2y = 4, now subtract (1/2)x from both sides getting -2y= (-1/2)x + 4 Now divide both sides by -2 getting: y = (1/4)x - 2 Do you agree that the first line is (in slope intercept form) y= (1/4)x -2?

OpenStudy (anonymous):

i think so

OpenStudy (radar):

Now lets take the second equation which is: 4y - x = -8 lets add x to both sides getting 4y = x - 8 Now divide thru by 4 getting: y=(1/4)x - 2 Do you see that they are identical?? One line is laying on top of the other, they are the same there is infinitely manty solutions.

OpenStudy (radar):

Can you use the "Draw" function and sketch your graph so we can look at it?

OpenStudy (radar):